OpenStudy (anonymous):

While traveling to Europe, Phelan exchanged 250 US dollars for euros. He spent 150 euros on his trip. After returning to the United States he converts his money back to US dollars. How much of the original 250 US dollars does Phelan now have? 1 European euro = 1.3687 US dollars 44.70 US dollars 73.06 US dollars 136.87 US dollars 140.41 US dollars

OpenStudy (texaschic101):

1/1.3687 = 150/x --- (1 euro to 1.3687 $ = 150 euro to x $) cross multiply x = (150)(1.3687) x = 205.305 250 - 205.305 = 44.695 riunds to 44.70